# Regularized second-order optimization of tensor-network Born machines

### v1

@@ Abstract
Tensor-network Born machines (TNBMs) are quantum-inspired generative models for learning data distributions. Using tensor-network contraction and optimization techniques, the model learns an efficient representation of the target distribution, capable of capturing complex correlations with a compact parameterization. Despite their promise, the optimization of TNBMs presents several challenges. A key bottleneck of TNBMs is the logarithmic nature of the loss function that is commonly used for this problem. The single-tensor logarithmic optimization problem cannot be solved analytically, necessitating an iterative approach that slows down convergence and increases the risk of getting trapped in one of many non-optimal local minima.
In this paper, we present an improved second-order optimization technique for TNBM training, which significantly enhances convergence rates and the quality of the optimized model. Our method employs a modified Newton’s method on the manifold of normalized states, incorporating regularization of the loss landscape to mitigate local minima issues. We demonstrate the effectiveness of our approach by training a one-dimensional matrix product state (MPS) on both discrete and continuous datasets, showcasing its advantages in terms of stability, efficiency, and generalization.
@@ I Introduction
The tensor-by-tensor scheme has an advantage for optimization of quadratic cost function, such as expectation values of observables, where single-tensor updates can be solved directly via eigenvalue problems. In these cases, each optimization step introduce a significant reduction of the cost function. However, for generative tasks using the negative log-likelihood (NLL) loss, the logarithmic nature of the landscape complicates optimization, requiring incremental steps of gradient-descent. This introduces two key challenges: first, the inner gradient-descent loop slows down training by adding a computational subroutine; second, unlike the analytical solutions in DMRG, numerical optimization is prone to nonoptimal local minima due to the abundance of singularities and minima in the NLL landscape.
@@ II.1 Tensor-network Born machines
Tensor-network Born machines describe distributions over bit strings using the amplitudes of a variational state, represented as a tensor network. Inspired by quantum physics, these amplitudes correspond to those of a quantum state and can take negative or complex values, while the probability of each bit string is given by the squared absolute value of its amplitude.
The quantum state itself is encoded as a tensor network, where a set of interconnected tensor cores represents its amplitudes. This structure is commonly visualized using a Penrose diagram [24], in which nodes correspond to tensor cores and edges indicate index contractions. Indices connecting two tensors are summed over, merging them into a single tensor, while free indices serve as input vertices.
To grade the performance of the TN, TNBM models use the negative log-likelihood (NLL) loss function. The NLL function is a measure of the between a probability distribution of the model $p(x)$ and a set of samples from the target distribution $\{x_{i}\}$, defined as
@@
where $p(x)$ is the likelihood of a sample predicted by the model. Up to a constant, the NLL function approximates the Kullback–Leibler divergence between the model and target distributions. For TNBMs, the predicted probability is given by Born rule $p(x)=\absolutevalue{\bra{\psi}\ket{x}}^{2}$, where |$\ket{\psi}$ and |$\ket{x}$ are the quantum state of the tensor network and the sample bitstring, respectively (see Fig. 1).
@@
In this work we focus on the one-dimensional tensor network variant known as the matrix product state (MPS), which is commonly used to model the ground state of 1D systems with gapped Hamiltonians [25, 26]. An MPS is a one-dimensional array of tensors, each having two bond indices connecting it to the tensors on the left and right, and a single free “site” index. At the edges of the system, we choose an open boundary condition, such that the first and last tensors have only a single bond index. The dimensions of the bond indices control the level of correlation of the probability distribution between different sites, quantified by the entanglement entropy of the quantum state.
@@
For the optimization process of the MPS, we look to update the different tensor cores of the state to minimize the loss function. There are several approaches to this task. A straightforward approach is to perform gradient descent on the different tensors simultaneously with a small learning rate. With this approach, in each iteration every tensor is updated slightly, often requiring hundreds or thousands of epochs converge. An alternative approach is to use local optimization scheme, where only a single tensor is updated at a time. Taking advantage of tensor contraction and simplification techniques, each single-tensor optimization problem is reformulated in terms of small tensors representing the environment induced by the tensor network. By sweeping across all the tensor cores sequentially, the compact representation can be retained and modified to describe the local optimization problem at each site, avoiding the need to repeatedly contract the entire tensor network. For MPS, this method forms the basis of the DMRG algorithm, widely used in condensed matter physics and quantum chemistry [21, 27, 28].
@@
For some choices of cost functions, the local optimization problems can be solved in a single step by analytically finding the minimal solution. This is the case for quadratic cost functions naturally used to describe the energy of a state in quantum systems, whose local optimization problem can be solved by singular value decomposition of the environment tensor [25]. For the case of the NLL cost function this is not the case, and the single tensor optimization tasks need to be solved iteratively using gradient-based optimization subroutine, slowing the optimization process. This computational bottleneck motivates the use of optimization techniques with faster convergence for the single-tensor sub-optimization problem. To address this, we study the option of performing Newton-method approach that operates on the constrained manifold of normalized tensors, improving convergence speed while maintaining numerical stability.
@@ II.2 Single tensor optimization
${L}(T)=-\sum_{x}{n_{x}\log\left(\frac{{(T,w_{x})}^{2}}{(T,T)}\right)},$ (4)
where $n_{x}$ is the frequency of the string $x$ within the selected samples, with $\sum_{x}{n_{x}}=1$. While the loss is defined for any choice of tensors, the physical interpretation of the MPS as a model of probability distribution constrains the space of valid tensors to the manifold of tensors with the constant norm of one- $(T,T)=1$.
Geometrically, this constraint manifold corresponds to a hypersphere $S^{D-1}$ of dimension $D-1$, where $D={\rm dim}(T)$. This puts our optimization problem within the framework of optimization on Riemannian manifolds
Following the notation used in [29], we use the notation $\overline{L}(T)$ for the unconstrained loss function, and $L(T)$ for the version constrained by the constant normalization condition.
@@ III.1 Second-order optimization algorithm
To optimize the NLL cost function using constrained Newton’s method, we first define gradients and Hessians in the context of constrained optimization. Unlike optimization of functions in free space, the derivatives of constraint functions include components that extend outside the constraint manifold. A simple gradient-based step would violate the constraint, requiring projection onto the tangent space. The gradients and Hessian tensors have to be adjusted to reside within the tangent space of the manifold, as described below.
@@
In this work, we apply Newton’s method to accelerate the optimization of each single tensor. Newton’s method refines the optimization step by incorporating second-order derivative information, adjusting both the step size and direction based on a local quadratic approximation of the loss function. Newton’s optimizer assumes that, near a minimum, the loss function can be well approximated by a quadratic model. Using the second derivative of $L(T)$, the update rule is given by:
$T\rightarrow T-\hat{H}^{-1}\nabla{L}$ (8)
where $H$ is the Hessian matrix and $\nabla L$ is the gradient vector of the loss function. In free space, the Hessian is defined as the matrix of second derivatives of the loss function with respect to pairs of variables
@@
${\rm Hess}\,\overline{L}(T)=2I+2\sum_{x}{n_{x}\frac{{w_{x}}{w_{x}}^{T}}{(T,w_{x})^{2}}}.$ (10)
For constrained optimization, Newton’s method has to be adapted to avoid large jumps out of the manifold. Unlike simple gradient descent, where the constraint could be accounted for by stepping in the direction of the gradient and projecting the new point back to the manifold, with Newton’s method this approach is ineffective. The step size can be quite large and can throw the optimizer far away from the original manifold. Moreover, it fails to take into account the curvature of the manifold itself, which is vital for second-order approximation of the loss function on the manifold.
@@
${\rm Hess}\,L(T)=2(I-T^{\otimes 2})+2\sum_{x}{n_{x}{\frac{\Pi_{T}(w_{x})^{\otimes 2}}{(T,w_{x})^{2}}}},$ (13)
where $\otimes$ denotes the outer product between two vectors. The expression above is obtained by taking the derivatives and projections in Eq. 12, and performing symmetrization on the resulting tensor, which simplifies the expression but has no effect on the action of the Hessian on vectors inside the tangent space.
@@
The first equation corrects the size and direction of the bare gradient while the second equation keeps the step within the tangent plane. Notice that the Hesian is singular by definition and is no longer invertible in $R^{D\times D}$.
@@ III.2 The local-minima problem
The constraint Newton’s optimization method has potential to accelerate the convergence of the algorithm as it contains a more accurate description of the loss function landscape. Nevertheless, using bare Newton steps as the sole mechanism of the optimization algorithm can lead to various problems, which are manifested strongly in the case of TNBMs optimization.
To understand the challenges of TNBM optimization, we examine a one-dimensional slice of the cost function, plotted in Fig. 2. As illustrated in Fig. 2(b), the landscape is filled with many asymptotic singularities, creating many narrow local minima. This structure persists in the multidimensional tensor space: the singularities appear across linear planes that dissect the landscape into many small regions, each with its distinct local minima. This is a natural property of the NLL loss function, as it is constructed from many logarithmic poles, one for each sample (see equation 4), and the overlap of the tensor $T$ with the different samples create singularity walls on the manifold that cannot be transversed using a continuous gradient flow. The logarithmic poles dissect the landscape into $2^{n_{s}}$ regions, corresponding to all possible configurations of overlap signs.
As a result, the type of the optimizer plays a key role for successful convergence. In steepest descent, the optimizer naturally avoids narrow minima by taking large steps and leaping away from narrow gorges with steep walls. This overshooting mechanism filters potential local minima with high second derivatives, and converges only at points that are relatively flat. In contrast, Newton’s method adjusts both the direction and magnitude of each step using the Hessian, and the optimizer tends to jump directly into one of the closest minimal points, adjusting the size of the step to avoid overshooting.
@@ III.3 Regularization of the NLL loss landscape
In the first regularization variant, we add a constant to the argument of each the logarithm, broadening and smoothing the singularities of the landscape
$\mathcal{L}_{\epsilon}(x)=\log\left(\absolutevalue{x}^{2}+\epsilon\right),$ (15)
where $\epsilon$ is a positive regularization constant. The sum of the broaden logarithms create the regularized loss function of the form
$L_{\epsilon}(\psi)=-\sum_{x}{n_{x}\mathcal{L}_{\epsilon}(\bra{\psi}\ket{x})}.$ (16)
@@
$\mathcal{L}_{\epsilon}(x)$ $=\mathcal{L}_{0}(x)*\kappa_{\epsilon}(x).$ (18)
The smoothing of the NLL function prevents sharp gradients and singularities on one hand, while producing an easy-to-calculate closed formula at the same time. Another important effect of regularization is a reduction in the number of local minima across the landscape. As shown in Fig. 2, increasing the regularization constant smooths out sharp features of the landscape, effectively eliminating more local minima as the constant grows. Meanwhile, the positions of the remaining minima remain largely stable, with only slight shifts in their positions.
Using the regularized loss function, we can apply Newton’s optimization method and avoid local minima while benefiting from the accelerated convergence of the second order optimization by Newton’s method.
The implementation of Newton’s method on the new regularized loss function requires caution, as several properties of the cost function have changed. Firstly, the projections of the gradient and Hessian have to be readjusted according to the new regularized loss function (see Appendix A). Secondly, the smoothing of the cost function replaces many singularities with local maxima points with concave regions around them. This affect negatively on the convergence of Newton’s optimizer, which rely on the convexity of the function, and can otherwise converge to any critical point, failing to distinguish between maxima and minima.
To optimize the non-convex cost function, we take an approach inspired by the trust region method [31, 32, 33, 34], which addresses this issue by bounding the size of the optimization step. In the trust-region method, the optimization step is determined using a quadratic approximation (or model) of the function, minimizing this model within a designated trust region— a region where the approximation is assumed to be valid. In cases when the model is convex and a minimum is available inside the trust region, the step coincides with the Newton step. Otherwise, when there is no minimal point within the interior of the defined trust region, the optimizer searches for a local minimum on the boundary of the region, shaped either as a hyper-sphere with a constant radius or, more generally, as a hyperboloid. By enforcing this constraint, the method benefits from second-order corrections even in non-convex regions while maintaining step-size control, ensuring stable optimization.
@@
In the case of the regularized NLL landscape, we can apply the principles of the trust region algorithm and correct the negative parts of the Hessian by adding quadratic terms as additional constraints. In our case, rather than adding a full scalar matrix and inhibiting the rate of convergence in all directions equally, we add terms that directly counter only the negative subspace of the Hessian. This can be done by taking the absolute values of all the terms that contribute to the negative part of the Hessian, which ensures that the modified matrix stays positive with minimal modification, which is equivalent to constraining the Newton step only in directions where the point is approaching a logarithmic singularity. Unlike the trust region method, we do not determine the size of the trust region apriori, rather we accept the step size resulted from this heuristic, although it is possible to expand our algorithm to include a precise control over the step size. The approach of adding additional terms as a numerical damping of the Hessian is also common in other numerical optimization algorithms, such as the Levenberg-Marquardt algorithm [35] and the Tikhonov regularization procedure in regression problems [36].
III.4 Other variants of regularization

### v2

@@ Abstract
Tensor-network Born machines (TNBMs) are quantum-inspired generative models for learning data distributions. Using tensor-network contraction and optimization techniques, the model learns an efficient representation of the target distribution, capable of capturing complex correlations with a compact parameterization. Despite their promise, the optimization of TNBMs presents several challenges. A key bottleneck of TNBMs is the logarithmic nature of the loss function commonly used for this problem. The single-tensor logarithmic optimization problem cannot be solved analytically, necessitating an iterative approach that slows down convergence and increases the risk of getting trapped in one of many non-optimal local minima.
In this paper, we present an improved second-order optimization technique for TNBM training, which significantly enhances convergence rates and the quality of the optimized model. Our method employs a modified Newton’s method on the manifold of normalized states, incorporating regularization of the loss landscape to mitigate local minima issues. We demonstrate the effectiveness of our approach by training a one-dimensional matrix product state (MPS) on both discrete and continuous datasets, showcasing its advantages in terms of stability and efficiency, and demonstrating its potential as a robust and scalable approach for optimizing quantum-inspired generative models.
@@ I Introduction
The tensor-by-tensor scheme has an advantage for optimization of quadratic cost function, such as expectation values of observables, where single-tensor updates can be solved directly via the eigenvalue problem. In these cases, each optimization step introduce a significant reduction of the cost function. However, for generative tasks using the negative log-likelihood (NLL) loss, the logarithmic nature of the landscape complicates optimization, requiring incremental steps of gradient descent. This introduces two key challenges: first, the inner gradient descent loop slows down training by adding a computational subroutine; second, unlike the analytical solutions in DMRG, numerical optimization is prone to nonoptimal convergence to local minima due to the abundance of singularities and minima in the NLL landscape.
@@ II.1 Tensor-network Born machines
Tensor-network Born machines describe distributions over bit-strings using the amplitudes of a variational state, represented as a tensor network. Inspired by quantum physics, these amplitudes correspond to those of a quantum state and can take negative or complex values, while the probability of each bit string is given by the squared absolute value of its amplitude.
Denoting the quantum state with the bra-ket notation as |$\ket{\psi}$, the predicted probability of the model is given by the Born rule $p(x)=\absolutevalue{\bra{\psi}\ket{x}}^{2}$, where |$\ket{\psi}$ is the inner quantum state of the Born machine and |$\ket{x}$ is the encoding of a sample bitstring as a quantum state.
To efficiently describe the state of the Born machine, the quantum state |$\ket{\psi}$ is represented as a tensor network, a network of interconnected tensor cores that reproduce global state when contracted together. This structure is commonly visualized using a Penrose diagram [24], in which nodes correspond to tensor cores and edges indicate index contractions. Indices connecting two tensors are summed over, merging them into a single tensor, while free indices serve as input vertices (See Fig. 1(a)).
To grade the performance of the TN, TNBM models use the negative log-likelihood (NLL) loss function. The NLL function measures how well the model’s probability distribution $p(x)$ matches the observed data $\{x_{i}\}$, defined as
@@
where $p(x)$ is the likelihood of a sample predicted by the model. Up to a constant, the NLL function approximates the Kullback–Leibler divergence between the model and target distributions.
@@
In this work we focus on the one-dimensional tensor network variant known as matrix product states (MPS), which are commonly used to model the state of 1D quantum systems [25, 26]. An MPS is a one-dimensional array of tensors, each having two bond indices connecting it to the left and right, and a single free “site” index representing the physical state. At the edges of the system, we choose an open boundary condition, such that the first and last tensors have only a single bond index. The dimensions of the bond indices control the level of correlation of the probability distribution between different sites, quantified by the entanglement entropy of the quantum state.
@@
For the optimization of the MPS, we look to update the different tensor cores of the state to minimize the loss function. There are several approaches to this task. A straightforward approach is to perform gradient descent on the different tensors simultaneously with a small learning rate. With this approach, in each iteration every tensor is updated slightly, often requiring hundreds or thousands of epochs to converge, which is often used for machine learning tasks. An alternative approach is to use local optimization scheme, where only a single tensor is updated at a time, especially when the cost function is quadratic, e.g. expectation value of a Hamiltonian. Taking advantage of tensor contraction and simplification techniques, each single-tensor optimization problem is reformulated in terms of the effective environment tensor defined as the contraction of the rest of the cores in the TN. By sweeping across all the tensor cores sequentially, the compact representation of the environment tensor can be retained and modified to describe the local optimization problem at each site (see Fig. 1(b)), avoiding the need to repeatedly contract the entire tensor network. For MPS, this method forms the basis of the DMRG algorithm, widely used in condensed matter physics and quantum chemistry [21, 27, 28].
@@
For some choices of cost functions, the local optimization problems can be solved in a single step by analytically finding the minimal solution. This is the case for quadratic cost functions naturally used to describe the energy of a state in quantum systems, whose local optimization problem can be solved by singular value decomposition of the environment tensor [25]. For the case of the NLL cost function this is not the case, and the single tensor optimization tasks need to be solved iteratively using gradient-based optimization subroutine, slowing the optimization process. This computational bottleneck motivates the use of optimization techniques with faster convergence for the single-tensor sub-optimization problem. To address this, we study the option of performing Newton-method optimization that operates on the constrained manifold of normalized tensors, improving convergence speed while maintaining numerical stability.
@@ II.2 Single tensor optimization
, where the last equal sign holds because the left and right environment tensors are isometries.
@@
${L}(T)=-\sum_{x}{n_{x}\log\left(\frac{{\left|(T,w_{x})\right|}^{2}}{(T,T)}\right)},$ (4)
where $n_{x}$ is the frequency of the string $x$ within the training set, with $\sum_{x}{n_{x}}=1$. While the loss is defined for any choice of tensors, the physical interpretation of the MPS as a model of probability distribution constrains the space of valid tensors to the manifold of tensors with the constant norm of one- $(T,T)=1$.
Geometrically, this constraint corresponds to the manifold of the hypersphere $S^{D-1}$ of dimension $D-1$, where $D={\rm dim}(T)=d\,\chi^{2}$, $d$ is the site index dimension and $\chi$ is the bond dimension of the MPS. This puts our optimization problem within the framework of optimization on Riemannian manifolds.
Following the notation used in [29], through this paper we use the notation $\overline{L}(T)$ for the unconstrained loss function, and $L(T)$ for the version constrained by the constant normalization condition.
@@ III.1 Second-order optimization algorithm
To optimize the NLL cost function using constrained Newton’s method, we first define gradients and Hessians in the context of constrained optimization. Unlike optimization of functions in free space, the derivatives of constraint functions include components that extend outside the constraint manifold. A simple gradient-based step could violate the constraint, requiring projection onto the tangent space. The gradients and Hessian tensors have to be adjusted to reside within the tangent space of the manifold, as described in the following.
@@
In this work, we apply Newton’s method to accelerate the optimization of each single tensor. Newton’s method refines the optimization step by incorporating second-order derivative information, adjusting both the step size and direction based on a local quadratic approximation of the loss function. This method has been applied to tensor network optimization in the cases of fidelity and expectation value optimization [30, 31, 32]. Newton’s optimizer assumes that, near a minimum, the loss function can be well approximated by a quadratic model. Using the second derivative of $L(T)$, the Newton step is give by
$T\rightarrow T-\left({\rm Hess}\,{L}\right)^{-1}{\rm grad}\,{L}$ (8)
where ${\rm Hess}\,{L}$ is the Hessian matrix and ${\rm grad}\,{L}$ is the gradient vector of the loss function. In free space, the Hessian is defined as the matrix of second derivatives of the loss function with respect to pairs of variables
@@
${\rm Hess}\,\overline{L}(T)=2I+2\sum_{x}{n_{x}\frac{{w_{x}}\otimes{w_{x}}}{(T,w_{x})^{2}}}.$ (10)
where $I$ is the identity matrix, and $\otimes$ denotes the outer product between two vectors.
For constrained optimization, Newton’s method has to be adapted to avoid large jumps out of the manifold. Unlike simple gradient descent, where the constraint could be accounted for by stepping in the direction of the gradient and projecting the new point back to the manifold, the step can be large and can throw the optimizer far away from the original manifold.
@@
${\rm Hess}\,L(T)=2(I-T^{\otimes 2})+2\sum_{x}{n_{x}{\frac{\left[\Pi_{T}(w_{x})\right]^{\otimes 2}}{(T,w_{x})^{2}}}}.$ (13)
The expression above is obtained by taking the derivatives and projections in Eq. 12, and performing symmetrization on the resulting tensor, which simplifies the expression but has no effect on the action of the Hessian on vectors inside the tangent space.
@@
The first equation corrects the size and direction of the bare gradient while the second equation keeps the step within the tangent plane. Notice that the Hesian is singular by definition and is no longer invertible in $\mathbb{R}^{D\times D}$.
@@ III.2 The local-minima problem
The constraint Newton’s optimization method has potential to accelerate the convergence of the algorithm as it contains a high-orderdescription of the loss function landscape. Nevertheless, using bare Newton steps as the sole mechanism of the optimization algorithm can lead to various problems, which are manifested strongly in the case of TNBMs optimization.
To understand these challenges, we examine a one-dimensional slice of the cost function, plotted in Fig. 2. As illustrated in Fig. 2(b), the landscape is filled with many asymptotic singularities, creating many narrow local minima. This structure persists in the multidimensional tensor space: the singularities appear across linear planes that dissect the landscape into many small regions, each with its distinct local minima. This is a natural property of the NLL loss function, as it is constructed from many logarithmic poles, one for each sample (see Eq. 4), and the overlap of the tensor $T$ with the different samples create singularity walls on the manifold that cannot be transversed using a continuous gradient flow. The logarithmic poles dissect the landscape into $2^{N_{s}}$ regions, corresponding to all possible configurations of overlap signs.
As a result, the choice of optimizer plays a key role in successful convergence. With steepest descent, the optimizer naturally avoids narrow minima by taking large steps and leaping away from narrow gorges with steep walls. This overshooting mechanism filters potential local minima with high second derivatives and converges only to points that are relatively flat. In contrast, Newton’s method adjusts both the direction and magnitude of each step using the Hessian, and the optimizer tends to jump directly into one of the closest minimal points, adjusting the size of the step to avoid overshooting.
@@ III.3 Regularization of the NLL loss landscape
In the first regularization variant, we add a constant to the argument of each logarithm, broadening and smoothing the singularities of the landscape. Given the regularized log function
$l_{\epsilon}(x)=\log\left(\absolutevalue{x}^{2}+\epsilon\right),$ (15)
where $\epsilon$ is a positive regularization constant, we define the regularized cost function as the sum of the broaden logarithms of the overlaps with
$L_{\epsilon}(\psi)=-\sum_{x}{n_{x}l_{\epsilon}(\bra{\psi}\ket{x})}.$ (16)
@@
$l_{\epsilon}(x)$ $=l_{0}(x)*\kappa_{\epsilon}(x).$ (18)
The smoothing of the NLL function prevents sharp gradients and singularities on one hand, while producing an easy-to-calculate closed formula at the same time. Another important effect of regularization is a reduction in the number of local minima across the landscape. As shown in Fig. 2(b), increasing the regularization constant smooths out sharp features of the landscape, effectively eliminating local minimum points as the constant grows. Meanwhile, the positions of the remaining minima remain largely stable, with no significant shifts in their positions.
Using the regularized loss function, we can reapply Newton’s optimization method and avoid local minima while benefiting from the accelerated convergence of the second order optimization.
The implementation of Newton’s method on the new regularized loss function requires caution, as several properties of the cost function have changed: the projections of the gradient and Hessian have to be readjusted according to the new regularized loss function (see Appendix A). In addition, the smoothing of the cost function replaces many singularities with local maxima points with concave regions around them (see Fig. 2(b)). This has a negative impact on the convergence of Newton’s optimizer, which rely on the convexity of the function, and can otherwise converge to any critical point, failing to distinguish between maxima and minima.
To address this, we take an approach inspired by the trust region method [34, 35, 36, 37], which mitigates the issue by bounding the size of the optimization step. In the trust-region method, the optimizer takes into account that the second order approximation is expected to be valid only within a bounded region of the search space (the trust region). As a result, the optimization step is bounded to either the interior of the region, if a local minimum exists within it, or to the boundary otherwise.
In cases the model is convex and a minimum is available inside the trust region, the step coincides with the Newton step. When there is no minimal point within the interior of the defined trust region, the optimizer searches for a local minimum on the boundary of the region, shaped either as a constant-radius hyper-sphere or, more generally, as a hyperboloid. By enforcing this constraint, the method ensures stable improvement at each step while benefiting from second-order corrections even in nonconvex regions.
@@
In the case of the regularized NLL landscape, we can apply the principles of the trust region algorithm and correct the negative parts of the Hessian by adding quadratic terms as additional constraints. In our case, rather than adding a full scalar matrix and inhibiting the rate of convergence in all directions equally, we add terms that directly counter only the negative subspace of the Hessian. This can be done by taking the absolute values of all the terms that contribute to the negative part of the Hessian, which ensures that the modified matrix stays positive with minimal modification (see Appendix A), which is equivalent to constraining the Newton step only in directions where the point is approaching a logarithmic singularity. Unlike the trust region method, we do not determine the size of the trust region apriori, rather we accept the step size resulted from this heuristic, although it is possible to expand our algorithm to include a precise control over the step size. The approach of adding additional terms as a numerical damping of the Hessian is also common in other numerical optimization algorithms, such as the Levenberg-Marquardt algorithm [38] and the Tikhonov regularization procedure in regression problems [39].
III.4 Second variant- regularization by shift